OpenStudy (anonymous):

I need some help with my trigonometry. I don't understand it. Could Somebody help?

OpenStudy (anonymous):

Give me a minute to post it.

OpenStudy (anonymous):

 Find a, b, and c

OpenStudy (anonymous):

@Godlovesme

OpenStudy (freckles):

You can find a by using Pythagorean theorem

OpenStudy (freckles):

then you can find one of the unknown angles by using a trig ratio

OpenStudy (freckles):

then you find the last angle by using that the sum of angle measurements of at triangle equal 180 deg

OpenStudy (anonymous):

Ok! thank you @freckles! I hate geometry, algebra is better.

OpenStudy (freckles):

well @GodGirl360 if you want to see if you can find the values yourself i will be here to check them but if you feel like you aren't sure what I'm talking about let me know we can try to work together

OpenStudy (anonymous):

@freckles a=12.84 b=146 degrees c=34 degrees

OpenStudy (freckles):

hmm... these things look a little off let's look at the hyp measurement we call a first

OpenStudy (freckles):

\[3.1^2+8.9^2=a^2 \\ \sqrt{3.1^2+8.9^2}=a\]

OpenStudy (freckles):

is this what you did for a?

OpenStudy (anonymous):

yes, i realized i did the math wrong while solving it. the answer is 9.42.

OpenStudy (freckles):

sounds great for a

OpenStudy (freckles):

now lets look at b

OpenStudy (freckles):

 I will round a to the nearest tenth since everything else is given in tenths but you can leave yours as hundredths unless your question says to round to hundredths or whatever